OpenStudy (anonymous):

Given f(x) = –5x + 1, what is f(–2)? A.-9 B.11 C.3/5 D.1

OpenStudy (opcode):

f(x) = –5x + 1, what is f(–2)? What is \(f(–2)\) basically just asks for you to sub in \(-2\) for \(x\) So: \[f(x) = –5x + 1 \implies f(-2) = -5(-2) + 1 \implies 10 + 1\] \[= 11\] So B is you answer. @kkbb240 if you do not understand just tag me and I will try to provide a better explanation. :-)

OpenStudy (anonymous):

Looks like @Opcode got there before i did :)

OpenStudy (anonymous):

thank u guys : )

OpenStudy (anonymous):

opcode can u help me with some more questions

OpenStudy (anonymous):

I can help you (:

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

Which is a solution to the equation 5x – y = –3? A. (2, –7) B. (1, 2) C. (–2, 7) D. (–2, –7)

OpenStudy (anonymous):

C

OpenStudy (anonymous):

What is the slope of the graph of 2x – 6y = 9? A -1/3 B -3 C 3 D 1/3

OpenStudy (anonymous):

6y=-2x+9 Simplify to: y=-1/3x+3/2 answer: A

OpenStudy (anonymous):

Which relation is a function?

OpenStudy (anonymous):

Graph B

OpenStudy (anonymous):

A carpenter charges a one-time fee of $50 plus $35 for each hour on the job. If n is the number of hours on the job and c is the total charge for the work done, then the variables are related by the equation c = 35n + 50. Which variable is the independent variable? A. n B. c C. There is no relationship between the variables.

OpenStudy (anonymous):

A

OpenStudy (anonymous):

What are the domain and range of the function {(3, 1), (5, –1), (9, –1), (2, 7), (–4, 1)}? A. Domain: {1, 7}; Range: {3, 5, 9, 2 } B. Domain: {1, –1, 7}; Range: {3, 5, 9, 2, –4} C. Domain: {3, 5, 9, 2, –4}; Range: {1, –1, 7} D. Domain: {3, 5, 9, 2 }; Range: {1, 7}

OpenStudy (anonymous):

C

OpenStudy (anonymous):

What are the intercepts of the line? A. x-intercept: 5; y-intercept: 4 B. x-intercept: –4; y-intercept: –5 C. x-intercept: –4; y-intercept: 5 D. x-intercept: 5; y-intercept: –4 http://static.k12.com/calms_media/media/1299500_1300000/1299540/1/c6db42a0fee13e426890e2d3f18de4f5c5146e83/143707.jpg

OpenStudy (anonymous):

D

OpenStudy (anonymous):

The cost of a product has been increasing each year. The graph shows data about the product’s cost over time. Which is a reasonable prediction for the cost of the product in year 12? A. $14 B. $10 C. $12 D. $18 http://static.k12.com/calms_media/media/245000_245500/245365/1/52f449f37f05b75c2160b6a2ad1163ebf9acdd1b/VHS_PA_S2_02_L214_L314-UT_Q10.gif